OpenStudy (anonymous):

Let and u=<4, -1> and v=<3,5>. What is |u+v| rounded to the nearest hundredth?

OpenStudy (anonymous):

Answer choices:

OpenStudy (reemii):

What do you have for |u+v| ?

OpenStudy (anonymous):

I'm not sure?

OpenStudy (reemii):

u+v = ?

OpenStudy (amistre64):

if you are not sure how to add vectors yet, then you are not prepared to solve this problem.

OpenStudy (amistre64):

in order to add vectors, you simply add like components <a,b> + <x,y> = <a+x, b+y> then the question is asking you for the length of the resultant vector ....

OpenStudy (anonymous):

I got <8, 4> by adding them..

OpenStudy (anonymous):

Now what?

OpenStudy (reemii):

you must apply the formula for the norm of a vector. what expression do you get?

OpenStudy (reemii):

norm of the vector \((a,b)\) is the quantity \(\sqrt{a^2+b^2}\).

OpenStudy (anonymous):

8^2 = 64 4^2 = 16 Square root 80 = 8.9

OpenStudy (reemii):

the sum is actually <7,4>. so the norm is sqrt(49+16) = sqrt(65) , that's almost sqrt(64), so ...